Reports 1-1 of 1 Clear search Modify search
Injection system (General activities)
Allocca, chiummo, de Rossi, Pillant - 23:35 Wednesday 23 January 2019 (44502) Print this report
Fast UnloCK investigation

Tonight a period of Fast UnloCK crisis troubled the injection system.

We exploited this pathological state to take data with different configurations:

 

At around 17h41 LT, we switched the IMC rampoto' roll-off for the piezo actuation from 1/f^4 to 1/f. The IMC immediately relocked and did not lose lock for some ten minutes until we switched back to 1/f^4 at 17h48.

At 20h13 LT, triggered by a new crisis, again we switched to 1/f. Same outcome.
20h23 LT back to 1/f^4


For the record, the other actions we tried:
20h30 LT 1/f
20h40 LT 1/f^4
(unlock NOT in coincidence with the switch, some 30 seconds afterwards...)


at 21h12 LT
IMC gain dB lock range: 10 (locks also at 7, but from refl image there is a fast oscillation)  to 22
at 21h14 set at 16 (was 15)


at 21h26 set back at 15dB


22h42 LT 1/f
(trying to lock arms - fast unlocks - set ITF_MAIN down)
(INJ_MAIN stuck in RFC_SCAN, ignores redirect)
(find RFC resonance by its own, then again FMODERR via MC_Z)

After recovering IMC_Restored, we tried again to lock the ITF, this time the IMC was much more stable. Lock acquisition attempts of the ITF were killed by other issues (see entry).

The rampoto' is left with 1/f configuration for the night. We will analyze data for the different periods to spot why with the 1/f the lock of the IMC seems much more robust even while trying to lock the ITF. This seems anyway to point towards something related to the EOM actuation.

Side note: no attempts to open the PStab loop or the alignment were made.
 

 

Comments to this report:
bersanetti - 2:22 Thursday 24 January 2019 (44505) Print this report

Until ~ 21:51 UTC everything seems fine, as INJ_MAIN was doing the flow FmodErr tuning with the loop (index 32, FmodErr_ACp going towards zero); then, during the scan of the RFC (index 34) the IMC started to unlock (with degraded power, alignment issue?), and the RFC ramp always starting in the same direction drove the MC mirror away. Once it was stable again, another FmodErr tuning was needed (two, actually) in order to revert this change in position of the MC mirror. Note that the ML thermal correction is reset at the end of the FmodErr procedure (index transition 32 -> 34) so that the good fringe of the RFC should always be close by.

Images attached to this comment
Search
×

Warning

×